OpenStudy (anonymous):

The system of equations is inconsistent. What are the missing values? y = 2/3x + 3 y = ( )/( )x + 5 help??

OpenStudy (anonymous):

\[y = \frac{ 2 }{ 3 }x + 3\]\[y = \frac{ ? }{ ? }x + 5\]

OpenStudy (anonymous):

@swissgirl @katti @dtan5457

OpenStudy (mathmate):

The only way two linear equations are inconsistent is they have the same slope but have different y-intercepts. In an equation y=ax+b, a is the slope, b is the y-intercept. If they had different slopes, they would eventually meet and give a unique solution.

OpenStudy (anonymous):

So how do I solve it? I have to fill in the blanks...how do you do that? @mathmate